OpenStudy (anonymous):

I need to factor b^3 - 8a^3. How would I start?

jimthompson5910 (jim_thompson5910):

notice how 8a^3 = (2a)^3

jimthompson5910 (jim_thompson5910):

so b^3 - 8a^3 is the same as b^3 - (2a)^3 from here you use the difference of cubes factoring rule to factor

jimthompson5910 (jim_thompson5910):

that rule is x^3 - y^3 = (x-y)(x^2+xy+y^2)